A merchant mixed 10 pounds of a cinnamon tea with 6 pounds of spice tea. The 16 pound mixture cost $35. A second mixture included 16 pounds of cinnamon tea and 4 pounds of spice tea. The 20 pound mixture cost $42. Find the cost per pound of the cinnamon tea and the spice tea.I came up with $2.14 for the cinnamon and $2.13 for the spice, but these answers were clearly wrong.
Answer by jorel1380(3719)

10c+6s=35
16c+4s=42
-----------------
48c+12s=126
20c+12s=70
28c=56
c=2
s=2.50
The cinnamon tea cost $2.00 a pound. The spice tea was $2.50 a pound..